15) Edgar and Cordelia meet in a UMUC Math 009 classroom and find that they have a great deal in common in addition to their great love of algebra. Among other things, they both love long-distance bicycling, so they decide to conduct an algebra experiment. Edgar, who lives in Richmond, Virginia, and Cordelia, who lives in Baltimore, Maryland, set out simultaneously on their bicycles on an agreed-upon route, cycling toward each other. If Edgar averages 12 mph and Cordelia averages 15 mph, and if the total distance between their starting points is 162 miles, how long will it take them to meet? (Round your answer to the nearest tenth of an hour).
Answer by mananth(16949)

Edgar averages 12 mph
Cordelia averages 15 mph,
total distance is 162 miles,
Let them meet at distance x from Edgar's starting point
Cordelia meets at a distance 162-x from her end.
The time they take to meet is same since both start at the same time.
..
time taken by Edgar = x/12
Time taken by Cordelia = 162-x/ 15
..
x/12 = 162-x / 15
15x=12(162-x)
15x= 1944-12x
15x+12x=1944
27x=1944
x= 1944/27
x=72 miles
..
distance traveled by Edgar = 72 miles
Edgar's average speed = 12 mph
time taken = 72/12
=6 hours
..
They meet at distance of 72 miles from Richmond and after 6 hours